# BAP v2: An Enhanced Task Framework for Instruction Following in Minecraft Dialogues

### v2

@@ 11 Appendix: Generating Synthetic Data
Sections 11.111.2, and 11.3 provide additional details (beyond those covered in Section 5) on how each of the three dialog simulators implements the general simulation framework (Section 5.1.2) – specifically, steps 1 (Architect Planning), 3 (Architect Instruction), and 4 (Optional Clarification Exchange). For details on the elementary shapes used for generating shape-based target structures (Section 5.3.1) see Section 11.4.
@@
All distances are measured using Manhattan Distance. For Steps 3 and 4, we outline only the key elements of our template-based generation framework. Further details will be included in our eventual code and data release. Thus, we omit detailed lexical and syntactic variations of utterances, providing only a few representative examples.
11.1 Random target structures and dialogs
Architect Planning
@@ 11.1 Random target structures and dialogs
The first four actions in the dialog are placements. Subsequent actions are sampled with a 90% probability for placement and 10% for removal. For placements, the algorithm selects a color uniformly randomly from the six colors and a location from a set of candidate positions. For removals, a block from the built structure is chosen at random. Candidate positions are defined as follows. If the board is empty (start of dialog), candidate positions are limited to the ground. But if at least one block exists, a candidate position must be within the "connected neighborhood" of an existing block. The "connected neighborhood" includes all locations within a Manhattan distance of 2 that share a face or edge with an existing block (but not a sole corner). This forms a 3x3x3 cube around the block, excluding the 8 diagonal corners but including the 6 adjacent and 12 2D diagonal locations.
Reference Block:
If needed, the algorithm selects a reference block as the closest existing block to the chosen next block. However, if the block from the last action that took place has the same distance, it takes precedence. A reference block can either be an existing block in the built structure (the last added one or a randomly selected one) or the last block that was removed from it. There are two special cases where a reference block is not required or cannot be assigned. These are discussed later in Section Special Cases with No Reference Block.
Architect Instruction
@@
The instruction specifies the color of the block, and when necessary, includes an optional adjective such as "floating" if temporary supporting blocks are required. For example, in Figure 11, "blue" and "floating" are used, while in Figure 12, "purple" and "yellow" are used.
@@
  1. 1.
    If the reference block is the last placed/removed block, the phrase used is "the last block you added/removed" etc. If the last action was a placement, the algorithm may omit the reference block in the instruction (ellipsis) with a small probability.
  2. 2.
    If the reference block’s color is unique in the built structure, it is referred to as "the [COLOR] block," where [COLOR] represents the block’s color.
  3. 3.
    If the color is not unique, the block is disambiguated using its spatial position relative to other blocks of the same color, allowing references such as "the leftmost [COLOR] block," etc.
For example, in Figure 11, the reference block is described as "the last block" and "the last block you placed." In Figure 12, it is referred to as "the closest yellow block" and "the orange one."
@@
The elements of the instruction that may be omitted include the location (i.e., the spatial relation + reference block) or the color of the next action. B’s clarification questions, such as "Where?" or "What color?" are then generated accordingly, followed by A’s response. For example, in Figure 11, the block’s color is omitted, prompting B to ask, "what color?" and A to reply, "red".
@@
If the next action is a removal and the block to be removed is the last placed block, the following adjustments are made while keeping everything else unchanged: 1) No reference block is needed. 2) A’s instruction will be a simple command such as "remove that block". 3) No clarification exchanges are generated, as there is nothing to omit in the instruction regarding the location or color of the block to be removed.
Start of Dialog
At the start of the dialog (empty board), the following modifications are made: 1. The first action is the of a block on the ground (randomly sampled from target structure blocks on the ground). 2. Since there are no prior blocks, A’s instruction will be a simple directive, such as "First start by placing an orange block on the ground".
11.2 Blocks-based dialogs for shape-based structures
Architect Planning
Recall that a target structure is provided as input to the dialog simulation. In this iteration of the simulation, blocks to be placed next are selected from those that remain unplaced, given the current built structure and the target. The algorithm first selects a single block to place. Based on this block and the reference block, it determines whether multiple blocks can be placed simultaneously while maintaining the same spatial relation. (We give an example of this later.) If so, multiple blocks are added; otherwise, only a single block is placed.
@@ 11.2 Blocks-based dialogs for shape-based structures
Next, we check whether multiple blocks (i.e., apart from the next block) can be placed "greedily" using the same reference block and spatial relation. For instance, if the next block is a yellow block placed to the left of the reference block, and another yellow block exists further to the left, A may plan to place both simultaneously for efficiency. To this end, the algorithm computes the direction of the next block relative to the reference block and checks if additional unplaced blocks of the same color exist along that direction. It then selects the longest contiguous sequence of such blocks connected to the next block.
Architect Instruction
The approach is similar to the one in Section 11.1 but additionally accounts for the number of blocks when multiple blocks are placed. For example, in Figure 16, the instruction specifies "three yellow blocks."
@@
There is only one special case, which is the same as the "Start of Dialog" one described in Section Special Cases with No Reference Block.
11.3 Shape-based dialogs for shape-based structures
Architect Planning
Recall that a target structure consisting of two shape instances is provided as input to the dialog simulation.
@@ 11.3 Shape-based dialogs for shape-based structures
Shapes are built sequentially. If both shape instances are on the ground, one is chosen randomly as the first to be built, otherwise the algorithm starts with the shape on the ground. Shapes are naturally built from the bottom up. The algorithm first selects a block within the shape to define spatial relations needed to specify where this shape should be built (e.g. wrt the reference block). This block also serves as the starting point for B when building the shape (the order of the remaining blocks is decided by B’s planning in Step 5 of the simulation algorithm of Section 5.1.2). Potential starting blocks are always located in the bottom corners of the shape. For the first shape, the starting block is the one farthest from the second shape. The starting block for the second shape is the block with the shortest distance to the last placed block in the first shape.
@@
Architect Instruction
@@
A shape’s size can be described in multiple ways depending on its dimensions. For 1D shapes such as rows, the description includes a single number, such as "four blocks long" or "three blocks tall." For 2D shapes like diagonals and planes, various formats can be used, such as "3x3", "two blocks long and three tall", etc. For example, in Figure 16, the instructions specify "4 yellow blocks" and "five block long."
The instruction also includes the shape’s color and name. We provide examples here for three shapes – rows, diagonals and planes. The primary shape names are "row," "diagonal," and "plane," though synonyms can be used. A horizontal row may be called a "line," while a vertical row may be referred to as a "column" or "tower." Similarly, a horizontal plane may be described as a "layer," whereas a vertical plane may be called a "wall." A horizontal diagonal is typically referred to as a "diagonal" or "diagonal line," but a vertical diagonal can also be described as a "staircase" or "stairway." For example, in Figure 16, the terms "pillar" and "line" are used for shape names, while "yellow" and "red" are used for shape colors.
@@
This applies only to the second shape being placed. Since the reference block is the last placed block, we use phrases such as "the last block you placed," similar to the approach in Sections 11.1 and 11.2.
@@
The approach is similar to the one in Sections 11.1 and 11.2, with two additional elements that can be omitted in an instruction. Recall that this data also includes spatial relations that use B as a spatial anchor (Section 5.1.3). For example, in Figure 16, A uses the spatial relation "going to the left of you" (using B as a spatial anchor) to indicate the direction for the row. Such directional spatial relations may be omitted from the instruction. Additionally, the dimensions/size of a shape may also be omitted.
11.4 Shape-based target structures
The six elementary shapes used to generate the structures are defined as follows:
@@ 11.4 Shape-based target structures
consist of at least 3 blocks, and are planar (along any of the $xy$, $yz$, or $xz$ planes). Horizontal diagonals lie along the $xz$ plane, vertical along either $xy$ or $yz$.
@@ 12 Appendix: Experimental Details
12.1 Curriculum Learning (CL) Algorithm
We use the Curriculum Arrangement algorithm from the CL framework proposed by Xu et al. (2020). Specifically, we sort the training dataset $D$ by the difficulty metric(s) and divided into $N$ equal-sized buckets $\{C_{i}:i=1,2,\dotsc,N\}$, corresponding to increasing difficulty levels from $C_{1}$ (easiest) to $C_{N}$ (hardest). The curriculum consists of $N$ stages $\{S_{i}:i=1,2,\dotsc,N\}$. In each stage $S_{i}$, a $1/N$ fraction of examples is sampled from each preceding bucket $\{C_{j}:j=1,2,\dotsc,i\}$, with the proportion of difficult examples gradually increasing across stages from 0 to eventually match the original dataset distribution. Within each stage, examples are still shuffled to maintain local stochastics, and there is no overlap between stages to prevent overfitting. At each stage, the model is trained for one epoch. After reaching $S_{N}$, the model should be ready for the original distribution in $D$. Hence, an additional stage $S_{N+1}$ is introduced, covering the entire dataset $D$, and training continues until convergence.
@@ 12.2 How much of each synthetic dataset to use during train time?
We provide further details of the procedure in Section 6.1. We first train models on the three synthetic datasets ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$) separately to study how increasing the training data affects performance. Focusing on the saturation regions of the learning curves—where performance plateaus and additional data has diminishing returns—we identify various "candidate" data amounts (#items) that yield near-peak performance for each dataset. Using these data amounts, we construct a grid of combinations across the three datasets, where each grid point corresponds to a different data mix and proportion of each dataset. For each grid point, we train a model on the combined $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$ and $D_{\mathrm{mc}}$ datasets. The combination that achieves the best performance on $D_{\mathrm{mc}}$ is selected. Thus, the approach yields a data mix that optimizes $D_{\mathrm{mc}}$ performance while also being highly performant on the synthetic data as much as possible. Also, in the best data mixes, the #items per dataset generally followed the order $D_{\mathrm{bs}}$ < $D_{\mathrm{ss}}$ < $D_{\mathrm{r}}$, which matches their difficulty ordering (as explored for CL in Section 6.2.1).
@@ 12.3 Experimental Setup for GRU-based Models
We train models using AdamW Loshchilov and Hutter (2019) and weight decay regularization with a weight decay factor of 0.1. The learning rate is set to 0.0001, and the batch size is 1. Early stopping is applied when the loss on the held-out development set increases monotonically for ten epochs. The development set for $M_{\mathrm{agg}}$ and $M_{\mathrm{agg+cl}}$, used for early stopping, is constructed from the validation splits of $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$ (Table 3) to match the dataset distribution in the training set.
@@
12.4 Experimental Setup for BERT-based Models
We provide further details for the experimental setup of $M_{\mathrm{mc}}^{\prime}$, $M_{\mathrm{agg}}^{\prime}$, and $M_{\mathrm{agg+cl}}^{\prime}$ (apart from those mentioned in Section 6.3.1).
Training
The training setup is similar to Section 12.3, with the following differences: a learning rate of 0.0001 for $M_{\mathrm{mc}}^{\prime}$ and 0.0005 for $M_{\mathrm{agg}}^{\prime}$ and $M_{\mathrm{agg+cl}}^{\prime}$, a batch size of 16, and a learning rate schedule consisting of linear warmup followed by inverse square root decay with 4000 warmup steps.
Model Hyperparameters
The transformer decoder consists of 5 layers with a representation dimensionality of 256, 2 attention heads, and an inner feedforward layer dimensionality of 512. The model uses uncased BERT with a maximum input length of 512 tokens, the highest allowed. Other hyperparameters are similar to Section 12.3.
13 Appendix: Additional Results
This section contains the BAP v2 metrics for the examples in Sections 7.2 and 8.1.
Example ID Model Type Color Location Shape Overall
Example 1 Baseline 1.0 1.0 0.0 0.0 0.0
Best 1.0 1.0 1.0 1.0 1.0
Example 2 Baseline 1.0 1.0 0.0 0.5 0.0
Best 1.0 1.0 1.0 1.0 1.0
Example 3 Baseline 1.0 1.0 0.0 0.0 0.0
Best 1.0 1.0 1.0 1.0 1.0
Table 10: BAP v2 metrics by Example ID for Section 7.2 (range $[0,1]$)
Example ID Model Type Color Location Shape Overall
Example A Baseline 0.67 0.67 0.0 0.0 0.0
Best 0.5 0.5 0.5 0.5 0.5
Example B Baseline 1.0 1.0 0.0 0.33 0.0
Best 1.0 1.0 0.0 1.0 0.0
Example C Baseline 0.5 0.5 0.0 0.0 0.0
Best 0.5 0.5 0.0 0.0 0.0
Example D Baseline 0.5 0.5 0.5 0.0 0.5
Best 1.0 1.0 0.5 0.0 0.5
Table 11: BAP v2 metrics by Example ID for Section 8.1 (range $[0,1]$)
@@ 14 Appendix: Annotation Setup
Our annotation team comprises one expert annotator (an author) and 10 non-expert annotators, who are undergraduate and graduate computer science students. The non-experts received elaborate training from the expert annotator before the final data collection. The NEB (non-empty board) subset is annotated by the 10 non-expert annotators, with each BAP item annotated by two annotators. Disagreements were resolved by the expert annotator, and the workload was evenly distributed among the annotators (in terms of the number of dialogs per annotator). The smaller and easier-to-annotate EB (empty board) subset was exclusively annotated by the expert annotator. Annotators were instructed to annotate one dialog at a time, processing items within a dialog in chronological order. This ensures they accurately understand the evolving game context and remain unbiased by future "unseen" events in a dialog.
15 Appendix: Extended Discussion on Concurrent Work
Apart from the discussion on concurrent work and Nebula’s evaluation in Section 8.3, a slightly more detailed comparison to Chaturvedi, Thompson, and Asher (2024) is provided in this section given some other relevant aspects of their work. (The detailed tables for Nebula’s performance on the full v2 benchmark mentioned in Section 8.3 are also provided here (Tables 13 and 14).)
15.1 Evaluation
Chaturvedi, Thompson, and Asher (2024) also identify some of the issues with BAP evaluation on $D_{\mathrm{mc}}$ that we do, and hence propose evaluation on synthetic data. They design simpler, non-dialog-based data with corresponding evaluation metrics. The data is single-turn and evaluates a model’s ability to follow an instruction in two scenarios: when the board is empty and the instruction is to build a single shape, or when a single shape is already present and the instruction involves placing or removing one block.
While this data is intentionally simple to test elementary Builder capabilities, it overly simplifies the task and therefore deviates significantly from the target MCBT/BAP. It lacks a dialog component, game history, and relative spatial references such as "left" and "right" as the Builder’s position and orientation are not considered. The world state/built structure is also fairly minimal, either empty or containing a single shape. While useful for elementary testing, it is a bit unclear how well insights from this data translate to the full BAP task. In contrast, our synthetic data remains simpler than $D_{\mathrm{mc}}$ while also closely emulating the MCBT, addressing these limitations and preserving many of the key BAP features and challenges, such as diverse instructions, rich spatial relations and floating blocks for example (more in Sections 2.1.1 and 5). Although designed/motivated for training more than evaluation, its simplicity allows assessment of basic model competencies, while its emulation of the MCBT enhances its applicability to the overall BAP task, thus striking a better balance from an evaluation standpoint. Section 8.2 details additional benefits of evaluating on our synthetic data.
Additionally, our synthetic data evaluation, combined with fairer F1, enables meaningful performance comparisons between synthetic data and $D_{\mathrm{mc}}$ (an example utility of this is the insights it provides to motivate Curriculum Learning). Chaturvedi, Thompson, and Asher (2024) are unable to conduct such comparisons (albeit understandably) due to the limitations of the legacy/original F1 metric.
15.2 Nebula+
While Chaturvedi, Thompson, and Asher (2024) primarily design their synthetic data for evaluation, they also report results from finetuning Nebula (Llama-3-8B) jointly on it and $D_{\mathrm{mc}}$. We refer to this model as Nebula+. (Recall that the original Nebula model was trained on $D_{\mathrm{mc}}$ alone.) Nebula+ improves performance on their synthetic data but remains similar to Nebula on $D_{\mathrm{mc}}$ (achieving an F1 score of 39.1% under the legacy F1, compared to Nebula’s 39.2%). This is fairly understandable given that their synthetic data, designed mainly for evaluation rather than training, does not capture sufficient task complexity, as discussed above. In contrast, as shown in Section 6, our approach of training jointly on our richer synthetic data ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$) and $D_{\mathrm{mc}}$ improves performance on both (and then further boosted by Curriculum Learning).
Similar to Nebula’s evaluation in Section 8.3, we evaluate Nebula+ here. Detailed results for the full v2 benchmark are provided in Tables 15 and 16. Table 12 presents Nebula and Nebula+ performance (micro F1) on both our synthetic data and the original $D_{\mathrm{mc}}$ test sets (extending the previous Table 9 with an additional row for Nebula+). The key trends and conclusions remain consistent with those previously discussed for Nebula—demonstrating the impact and robustness of our v2 evaluation, providing a fairer model assessment, and reinforcing the importance of evaluating on our synthetic data. Additionally, Table 12 shows that Nebula+ slightly underperforms compared to Nebula on our synthetic data while maintaining similar performance on $D_{\mathrm{mc}}$. This suggests that improvements on their synthetic data do not necessarily translate to better performance on ours (fairly understandable as their synthetic data does not capture sufficient task complexity).
Model Test Data
$D_{\mathrm{bs}}$ $D_{\mathrm{ss}}$ $D_{\mathrm{r}}$ $D_{\mathrm{mc}}$
Nebula 51.5 33.1 30.9 50.1
Nebula+ 48.8 32.4 30.4 50.1
Table 12: Nebula and Nebula+’s performance across datasets (v2 evaluation on $D_{\mathrm{mc}}$)
Dataset Type Color Location Shape Overall
EB 92.8 92.8 78.3 84.8 78.3
NEB 79.0 77.4 47.2 59.0 46.4
Overall 80.6 79.1 50.7 61.9 50.1
Table 13: Micro F1 scores for the Nebula model under the v2 benchmark
Dataset Type Color Location Shape Overall
EB 96.8 96.8 84.5 90.5 84.5
NEB 81.7 80.9 47.1 64.1 46.5
Overall 82.9 82.2 50.1 66.3 49.6
Table 14: Macro F1 scores for the Nebula model under the v2 benchmark
Dataset Type Color Location Shape Overall
EB 94.8 94.8 80.6 83.7 80.6
NEB 78.9 77.7 46.9 58.7 46.3
Overall 80.7 79.6 50.7 61.6 50.1
Table 15: Micro F1 scores for the Nebula+ model under the v2 benchmark
Dataset Type Color Location Shape Overall
EB 97.5 97.5 83.6 88.9 83.6
NEB 82.8 82.0 48.6 64.1 48.1
Overall 84.0 83.3 51.5 66.1 51.0
Table 16: Macro F1 scores for the Nebula+ model under the v2 benchmark

### v3

@@ 10 Appendix: Generating Synthetic Data
    Sections 10.210.3, and 10.4 provide additional details (beyond those covered in Section 5) on how each of the three dialogue simulators implements the general simulation framework (Section 5) – specifically,
    Step 10.1.
    1 (Architect Planning),
    Step 10.2.
    3 (Architect Instruction), and
    Step 10.3.
    4 (Optional Clarification Exchange). Section 10.5 provides details on the elementary shapes used to define shape-based target structures (Section 5.4.1).
@@
    All distances are measured using Manhattan Distance. For
    Step 10.4.
    3 and
    Step 10.5.
    4, we outline only the key elements of our template-based generation framework. Further details will be included in our eventual code and data release. Thus, we omit detailed lexical and syntactic variations of utterances, providing only a few representative examples.
    10.1 Target structures
    Some global constraints on target structures are enforced during their generation:
    1. (a)
      The structure must have at least one block on the ground.
    2. (b)
    10.2 Random target structures and dialogues ($D_{\mathrm{r}}$)
    Step 10.6.
    1: Architect Planning
@@ 10.2 Random target structures and dialogues ($D_{\mathrm{r}}$)
    The first four actions in the dialogue are placements. Subsequent actions are sampled with a 90% probability for placement and 10% for removal. For placements, the algorithm selects a color uniformly randomly from the six colors and a location from a set of candidate positions. For removals, a block from the built structure is chosen at random. Candidate positions are defined as follows. If the board is empty (start of dialogue), candidate positions are limited to the ground. But if at least one block exists, a candidate position must be within the ”connected neighborhood” of an existing block. The ”connected neighborhood” includes all locations within a Manhattan distance of 2 that share a face or edge with an existing block (but not a sole corner). This forms a 3x3x3 cube around the block, excluding the 8 diagonal corners but including the 6 adjacent and 12 2D diagonal locations.
    Step 10.8.
    3: Generating the Main Dialogue
@@
    The instruction specifies the color of the block (e.g. “blue”), and when necessary, includes an optional adjective such as “floating” if temporary supporting blocks are required.
@@
    1. (a)
      If the reference block is the last placed/removed block, the phrase used is “the last block you added/removed” etc. If the last action was a placement, the algorithm may omit the reference block in the instruction (ellipsis) with a small probability.
    2. (b)
      If the reference block’s color is unique in the built structure, it is referred to as “the [COLOR] block,” where [COLOR] represents the block’s color.
    3. (c)
      If the color is not unique, the block is disambiguated using its spatial position relative to other blocks of the same color, allowing references such as “the closest/leftmost [COLOR] block,” etc.
@@
    The elements of the instruction that may be omitted include the location (i.e., the spatial relation + reference block) or the color of the next action. B’s clarification questions, such as “Where?” or “What color?” are then generated accordingly, followed by A’s response.
@@
    If the next action is a removal and the block to be removed is the last placed block, the following adjustments are made while keeping everything else unchanged: 1) No reference block is needed. 2) A’s instruction will be a simple command such as “remove that block”. 3) No clarification exchanges are generated, as there is nothing to omit in the instruction regarding the location or color of the block to be removed.
    Start of Dialogue
    At the start of the dialogue (empty board), the following modifications are made: (a) The first action is the of a block on the ground (randomly sampled from target structure blocks on the ground). (b) Since there are no prior blocks, A’s instruction will be a simple directive, such as “First start by placing an orange block on the ground”.
    10.3 Blocks-based dialogues for shape-based structures ($D_{\mathrm{sb}}$)
    Step 10.10.
    1: Architect Planning
    Recall that a target structure is provided as input to the dialogue simulation. In this iteration of the simulation, blocks to be placed next are selected from those that remain unplaced, given the current built structure and the target. The algorithm first selects a single block to place. Based on this block and the reference block, it determines whether multiple blocks can be placed simultaneously while maintaining the same spatial relation. (We give an example of this later.) If so, multiple blocks are added; otherwise, only a single block is placed.
@@ 10.3 Blocks-based dialogues for shape-based structures ($D_{\mathrm{sb}}$)
    Next, we check whether multiple blocks (i.e., apart from the next block) can be placed ”greedily” using the same reference block and spatial relation. For instance, if the next block is a yellow block placed to the left of the reference block, and another yellow block exists further to the left, A may plan to place both simultaneously for efficiency. To this end, the algorithm computes the direction of the next block relative to the reference block and checks if additional unplaced blocks of the same color exist along that direction. It then selects the longest contiguous sequence of such blocks connected to the next block.
    Step 10.12.
    3: Generating the Main Dialogue
    The approach is similar to the one in Section 10.2 but additionally accounts for the number of blocks when multiple blocks are placed (e.g. to specify “three yellow blocks”).
@@
    There is only one special case, which is the same as the ”Start of Dialogue” one described in Section Special Cases with No Reference Block.
    10.4 Shape-based dialogues for shape-based structures ($D_{\mathrm{ss}}$)
    Step 10.14.
    1: Architect Planning
    Recall that a target structure consisting of two shape instances is provided as input to the dialogue simulation.
@@ 10.4 Shape-based dialogues for shape-based structures ($D_{\mathrm{ss}}$)
    Shapes are built sequentially. If both shape instances are on the ground, one is chosen randomly as the first to be built, otherwise the algorithm starts with the shape on the ground. Shapes are naturally built from the bottom up. The algorithm first selects a block within the shape to define spatial relations needed to specify where this shape should be built (e.g. wrt. the reference block). This block also serves as the starting point for B when building the shape (the order of the remaining blocks is decided by B’s planning in Step 5 of the simulation algorithm of Section 5). Potential starting blocks are always located in the bottom corners of the shape. For the first shape, the starting block is the one farthest from the second shape. The starting block for the second shape is the block with the shortest distance to the last placed block in the first shape.
@@
    Step 10.16.
    3: Generating the Main Dialogue
@@
    A shape’s size can be described in multiple ways depending on its dimensions. For 1D shapes such as rows, the description includes a single number, such as “four blocks long” or “three blocks tall.” For 2D shapes like diagonals and planes, various formats can be used, such as “3x3”, “two blocks long and three tall”, etc.
    The instruction also includes the shape’s color and name. We provide examples here for three shapes – rows, diagonals and planes. The primary shape names are “row,” “diagonal,” and “plane,” though synonyms can be used. A horizontal row may be called a “line,” while a vertical row may be referred to as a “column”, “pillar”, or “tower.” Similarly, a horizontal plane may be described as a “layer,” whereas a vertical plane may be called a “wall.” A horizontal diagonal is typically referred to as a “diagonal” or “diagonal line,” but a vertical diagonal can also be described as a “staircase” or “stairway.”
@@
    This applies only to the second shape being placed. Since the reference block is the last placed block, we use phrases such as “the last block you placed,” similar to the approach in Sections 10.2 and 10.3.
@@
    The approach is similar to the one in Sections 10.2 and 10.3, but now, the size/dimension and relative location/orientation of the shape to be placed (using expressions such as “going to the left of you” that use B as a spatial anchor, c.f. Section 5.2.3) can also be omitted.
    10.5 Elementary shapes
    The six elementary shapes used to generate the structures for $D_{\mathrm{sb}}$ and $D_{\mathrm{ss}}$ are defined as follows:
@@ 10.5 Elementary shapes
    consist of at least 3 blocks, and are planar (along any of the $xy$, yzyz, or $xz$ planes). Horizontal diagonals lie along the $xz$ plane, vertical along either $xy$ or $yz$.
@@ 11.1 How much of each synthetic dataset to use during train time?
    We provide further details of the procedure in Section 6.1. We first train models on the three synthetic datasets ($D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$) separately to study how increasing the training data affects performance. Focusing on the saturation regions of the learning curves—where performance plateaus and additional data has diminishing returns—we identify various ”candidate” data amounts (#items) that yield near-peak performance for each dataset. Using these data amounts, we construct a grid of combinations across the three datasets, where each grid point corresponds to a different data mix and proportion of each dataset. For each grid point, we train a model on the combined $D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$ and $D_{\mathrm{mc}}$ datasets. The combination that achieves the best performance on $D_{\mathrm{mc}}$ is selected. Thus, the approach yields a data mix that optimizes $D_{\mathrm{mc}}$ performance while also being highly performant on the synthetic data as much as possible. Also, in the best data mixes, the #items per dataset generally followed the order $D_{\mathrm{sb}}$ ¡ $D_{\mathrm{ss}}$ ¡ $D_{\mathrm{r}}$, which matches their difficulty ordering.
@@ 11.2 Experimental Setup for GRU-based Models
    We train models using AdamW Loshchilov and Hutter (2019) and weight decay regularization with a weight decay factor of 0.1. The learning rate is set to 0.0001, and the batch size is 1. Early stopping is applied when the loss on the held-out development set increases monotonically for ten epochs. The development set for $M_{\mathrm{agg}}$ and $M_{\mathrm{agg+cl}}$, used for early stopping, is constructed from the validation splits of $D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$ (Table 3) to match the dataset distribution in the training set.
@@ 12 Appendix: Annotation Setup
    Our annotation team comprises one expert annotator (an author) and 10 non-expert annotators, who are undergraduate and graduate computer science students. The non-experts received elaborate training from the expert annotator before the final data collection. The NEB (non-empty board) subset is annotated by the 10 non-expert annotators, with each BAP item annotated by two annotators. Disagreements were resolved by the expert annotator, and the workload was evenly distributed among the annotators (in terms of the number of dialogues per annotator). The smaller and easier-to-annotate EB (empty board) subset was exclusively annotated by the expert annotator. Annotators were instructed to annotate one dialogue at a time, processing items within a dialogue in chronological order. This ensures they accurately understand the evolving game context and remain unbiased by future ”unseen” events in a dialogue.